# Reasoning Abilities of Large Language Models: In-Depth Analysis on the Abstraction and Reasoning Corpus

### v1

@@ 3.2.4. Conclusion
In Section 3.2, we conducted experiments to evaluate the compositionality of LLMs in solving ARC tasks using DSLs. The results lead to the following two conclusions. First, LLM contains an ability to understand inputs including DSLs and grids. Second, the ability to analyze the combination of steps to decompose the task into smaller sub-tasks and achieve the desired result is weak. This indicates that while LLMs have a good understanding ability, they cannot still perform step-by-step planning.
The result of the experiment with a specific DSL for addressing ARC tasks also shows that the compositionality of LLMs is insufficient. Over the DSL used in this experiment, there are other DSLs designed to evaluate the compositional capabilities of LLMs such as Hodel (Hodel, 2023b) and Wind (a.k.a. icecuber) (Wind, 2020). However, even if LLMs could fully understand and use these more detailed DSLs, LLMs will be evaluated as lacking in compositional abilities due to the described limitations(the tendency to repeat the same DSL, the inability to find the right combination of DSLs for the answer). Furthermore, if the compositional abilities of the current LLM validator, which evaluates steps, are found lacking, then for LLMs to autonomously combine DSLs to solve ARC problems would be extremely challenging.
@@ 3.3. Capability of LLMs 3: Productivity
In Section 3.3, we investigate the third concept of LoTH, productivity. Productivity refers to the ability to generate unseen representations based on observed data. This characteristic allows humans to imagine different situations even from a single phenomenon, thus allowing for efficient learning without the need to learn from data inefficiently each time. Similarly, when equipped with this ability, LLMs are expected to excel in unseen tasks, making it one of the key functions of essential reasoning. The ability to generate new pairs within a limited set of rules could be helpful while solving ARC tasks, highlighting the need for productivity. In this section, we will test productivity by evaluating the validity of created examples based on given example pairs of ARC tasks.
3.3.1. Experiment
We conduct experiments using ARC tasks to understand how well LLMs can generate new expressions based on inherent logical concepts. Productivity involves two main steps: inferring specific rules for generating images from example images and natural language expressions and using those rules to generate new, unseen images. However, solving ARC tasks, experimented on in the previous sections so far, is unsuitable for confirming these two processes. For precise evaluation, we propose a new experiment: Given an ARC task and a basic rule shared with similar ARC tasks, can LLMs generate valid examples of the given task? If LLMs can understand a relationship between the given ARC task and the abstract rule, they should be able to infer specific rules for the given task and generate new valid examples. Through this, we want to see if LLMs can imitate human thinking productivity.
3.3.2. Method
@@
We proposed the Inverse Transformation Prompt (ITP), a prompting technique for this experiment. ITP instructs LLMs to generate multiple valid examples using the ARC task and its abstract rules. Fig. 15 demonstrates the process by which LLMs generate new examples given the ARC task and the corresponding ITP. LLMs generate multiple inputs that can form pairs with the output from one example of the task. This example, as mentioned earlier, is selected to serve as the basis for generation and is not included in ITP. If LLMs understood the specific rules corresponding to the ARC task given through ITP, the new example pairs generated by LLMs would be suitable as examples of the task.
@@
Before analyzing the experimental results, it was necessary to redefine the evaluation metric as the object was changed from solving ARC tasks to generating valid examples. As explained earlier, for one example of a particular ARC task, we generated valid inputs that could be paired with the output of that example. To properly generate inputs, LLM must understand the specific rules of the given ARC task through its ITP and apply the understood rules to the output for generating valid inputs. In this experiment, we evaluated whether all generated inputs were valid for each ARC task. This evaluation metric can check both whether the LLM understands the correct rules and whether it generates valid examples based on the rules it understands. Therefore, this experiment systematically evaluates the logical and valid demo pair generation capability of LLMs, contributing to our understanding of their ability to generate new representations.
3.3.3. Results
Based on 160 ARC tasks classified by ConceptARC, we evaluated the validity of a total of 2,913 generated examples. The valid generation ratio on average was approximately 17.1%, while the rest were invalid. As we mentioned before, the validity of results was determined by human judgment regarding whether the generated task adhered to the given rule. Results in Table 4 indicate that LLMs demonstrate a certain level of performance in generating examples consistent with the rule. However, since the criteria for validating the generated results as valid or invalid are weak, there is a limitation that results cannot be used before data post-processing even if an infinite number of results can be created. We analyzed the generated inputs to investigate why LLMs were unable to generate valid inputs and identified the following two sources of error: LLMs struggled with rule inference and were weak in step-by-step generation.
@@ 3.3.3. Results
Weak rule inference
Direct replication of other grids in the same task was a recurring issue, even with prompts designed to prevent such outcomes. As shown in Fig. 17, LLMs often simply copy the input from given example pairs to create new inputs. This error stemmed from LLMs’ failure to deduce meaningful rules from provided examples.
Weak step-wise productivity
There was another limitation in considering small steps to generate inputs from the given output. As a result, it was challenging to determine whether generated examples could be solved by specific rules of the task, leading to frequent occurrences of generated examples that could not be solved by rules of the task. Fig. 17 is the example. In the case of an input where all the vertices of the square have been erased, the color of the vertices cannot be determined, so the given output cannot be inferred.
@@ 3.3.4. Conclusion
In Section 3.3, we conducted experiments to confirm the productivity of LLMs by assessing whether they can understand given ARC tasks in abstracted representations and generate valid new examples based on abstracted rules. Although it is known that LLMs have great strengths in creating creative works, our experimental results reveal that LLMs are weak in understanding rules and producing creations that adhere to those rules. These results indicate that the process by which LLMs generate outputs is closer to mimicking human-generated results and achieving human-level generation abilities for LLMs is challenging.
Further research is needed to understand the specific mechanisms behind LLMs’ generation and how to elevate it to the human level. The experiments proposed in this study only examined whether LLMs can generate the given ARC tasks with simple rules and examples. However, it is difficult to determine whether the generated tasks were created following a human-like generation process or if they simply appear valid. Similarly, there is a need to analyze whether the process of incorrect generation resembles that of humans or not.
@@ 4.1.1. Abstract Knowledge
To solve ARC, the first challenge is the ability to extract the implicit information contained within ARC. Xu et al. argued that object-based representation is crucial for solving ARC and proposed ARGA (Xu et al., 2023a), which converts given example grids into graphs. Their follow-up study (Xu et al., 2023b) involved LLM solving ARC tasks using information obtained from ARGA and showed notable performance for object-based ARC tasks. However, these studies have a fundamental weakness in that they cannot be applied to ARC tasks without objects. Since only about 40% of ARC tasks contain object concepts (Xu et al., 2023a), this method cannot be applied to more than half of the tasks. Wang et al. on the other hand, improved the abstraction ability of LLMs to some extent with a graph-form dataset consisting of 221K textual descriptions, called AbsPyramid (Wang et al., 2023a), and also proposed a framework called AbsInstruct (Wang et al., 2024a) utilizing this dataset. Attempting to structure sentences can be an effective abstraction method for natural language, but its effectiveness cannot be seen in tasks that do not contain sentences.
@@ 4.1.2. Reasoning
Another challenge for LLMs in the context of ARC is the vast search space. One method gaining attention to address this is to enable LLMs to generate DSLs themselves. Rajani et al. introduced CAGE (Rajani et al., 2019), which prompts LLMs to generate explanations before generating answers. Subsequently, Wang et al. (Wang et al., 2024b) reported improved results by having LLMs generate DSLs based on hypotheses they set themselves. Additionally, active research is underway on prompting techniques applying algorithmic approaches. Zhou et al. (Zhou et al., 2022) demonstrated enhanced inference performance in LLMs by applying in-context learning. Follow-up research is actively being conducted following CoT and ToT. For example, CoT-SC (Wang et al., 2023b) is a study that selects results through voting from multiple instances of CoT, GoT (Besta et al., 2023) secures flexibility by enabling the generation of graph-like thought nodes, and XoT (Ding et al., 2023) uses the thought tree while Monte Carlo tree search and refines the tree with reinforcement learning. However, these attempts are closer to additional learning for LLMs, and more researches are needed to ascertain whether fundamental improvements in LLMs’ reasoning abilities are achievable.
@@ 5. Conclusions
In this study, we addressed the limitation of existing research, which predominantly analyzed LLM’s inference ability from a deterministic perspective, by introducing the LoTH perspective to ensure a fair evaluation of the process as well. By comparing and analyzing arguments across various domains of inference, we confirmed that logicality, compositionality, and productivity are quantifiable components to evaluate inference ability. Next, we proposed three experiments using the ARC dataset to quantitatively evaluate these three components from the LoTH perspective. The results showed that although current LLMs exhibit outstanding performance, they lack logicality, compositionality, and productivity in their processes, suggesting that they are closer to probabilistic mimicry rather than possessing autonomous inference abilities. Finally, we explored meaningful research directions for LLM to acquire inference capabilities from the LoTH perspective, as well as alternative approaches beyond ARC. This attempt to quantitatively evaluate the inference process from the LoTH perspective through experimental methods represents a differentiated approach not present in previous research. It contributes by providing a new perspective on how to handle inference capabilities in the field of computer science, going beyond LLMs.
@@ Acknowledgements.
This work was supported by the IITP (RS-2023-00216011) and the National Research Foundation (RS-2023-00240062) grants funded by the Ministry of Science and ICT, Korea. Experiments are supported by the Accelerate Foundation Models Research Initiative, Microsoft.References
@@
- Hodel (2023a) Michael Hodel. 2023a. ARC: Where Do We Stand Today? https://lab42.global/arc-article/
- Hodel (2023b) Michael Hodel. 2023b. Domain Specific Language for the Abstraction and Reasoning Corpus. https://github.com/michaelhodel/arc-dsl
- Huang and Chang (2023) Jie Huang and Kevin Chen-Chuan Chang. 2023. Towards Reasoning in Large Language Models: A Survey. In ACL.
@@
- Lab42 (2024) Lab42. 2024. ARCathon Leaderboard. https://lab42.global/arcathon/leaderboard/
@@
- Wang et al. (2023b) Xuezhi Wang, Jason Wei, Dale Schuurmans, Quoc Le, Ed Chi, Sharan Narang, Aakanksha Chowdhery, and Denny Zhou. 2023b. Self-Consistency Improves Chain of Thought Reasoning in Language Models. In ICLR.
@@
- Wang et al. (2023a) Zhaowei Wang, Haochen Shi, Weiqi Wang, Tianqing Fang, Hongming Zhang, Sehyun Choi, Xin Liu, and Yangqiu Song. 2023a. AbsPyramid: Benchmarking the Abstraction Ability of Language Models with a Unified Entailment Graph. arXiv:2311.09174 (2023).

### v2

@@ 3.2.5. Conclusion
In Section 3.2, experiments using ARC and DSL were conducted to measure the compositionality of LLMs. The results led to three conclusions. First, LLMs were able to predict the output grid when DSL was applied to the input with an average accuracy of about 81%. However, as the length of steps increased, the prediction rate decreased, which appears to be due to cumulative errors. Second, when not given the correct answer, LLMs selected the correct DSL only 3% of the time, indicating a lack of ability both in inferring rules to predict the correct output grid and in selecting the appropriate DSL to reach the expected output. Finally, when human descriptions were added, the accuracy in choosing DSL increased to a level similar to when the correct answer was provided. Analysis of this process suggested that this improvement was due to linguistic abstraction of the ARC task and DSL combinations.
Previous studies have emphasized LLMs’ limitations in combining simple elements to create new meanings, revealing struggles with compositionality. One study shows Transformers exhibit significant performance drops when tested on new function combinations, indicating challenges in systematically generalizing knowledge (Hupkes et al., 2020). Another study introduced datasets like SADE to evaluate LLMs’ ability to process visual and textual information, suggesting they still struggle with tasks like understanding negations and grasping complex content (Ma et al., 2024). A further study examined how well LLMs can break down complex instructions or build them from simple ones. These found that while LLMs improve at understanding simple tasks by learning complex ones, they struggle with complex tasks when starting from simpler ones (Yang et al., 2024). These findings across studies point to ongoing challenges in LLMs’ ability to connect simple and complex elements, highlighting their compositionality limitations.
@@ 3.3. Capability of LLMs 3: Productivity
3.3.1. Motivation
In Section 3.3, we investigate the third concept of the Language of Thought Hypothesis (LoTH): productivity. Productivity refers to the ability to generate unseen representations based on observed data (Fodor and Pylyshyn, 1988). This characteristic enables humans to imagine diverse situations from a single phenomenon, facilitating efficient learning without the need for repetitive data exposure. Similarly, when endowed with this ability, LLMs are expected to excel in unseen tasks, making productivity a crucial function of essential reasoning. The capacity to generate new pairs within a constrained set of rules is particularly valuable for solving ARC tasks, highlighting the need for productivity. In this section, we will assess productivity by evaluating the validity of LLM-generated examples based on given example pairs from ARC tasks.
While productivity ideally involves testing for infinite generative capacity, practical limitations necessitate alternative approaches. The challenge lies in demonstrating that a system can produce an unlimited number of novel, meaningful outputs from a finite set of inputs and rules. Previous studies have addressed this challenge by examining whether valid outputs can be produced under added constraints (van der Velde et al., 2004; Lake and Baroni, 2018; Hupkes et al., 2020). These constraints serve to create a more manageable testing environment while still allowing for the assessment of generative capabilities. Following this methodology, our study investigates how effectively LLMs can generate valid outputs when presented with an ARC task and its underlying conceptual rule. This approach allows us to evaluate productivity within a controlled framework while still capturing the essence of generative capacity.
To understand how well LLMs can generate new expressions based on inherent logical concepts, we conduct experiments using ARC tasks. Productivity in this context involves two main steps: 1) inferring specific rules for image generation from example images and natural language expressions, and 2) applying these rules to generate new, unseen images. However, the standard approach to solving ARC tasks, as explored in previous sections, is insufficient to confirm these two processes. Therefore, we propose a novel experiment: Given an ARC task and a basic rule shared with similar ARC tasks, can LLMs generate valid examples of the given task? If LLMs can understand the relationship between the given ARC task and the abstract rule, they should be able to infer specific rules for the task and generate new valid examples. Through this, we aim to determine whether LLMs can imitate the productivity of human thinking in generating novel solutions.
3.3.2. Validity of Augmentation
@@
We proposed the ITP, a prompting technique for this experiment. ITP instructs LLMs to generate multiple valid examples using the ARC task and its abstract rules. Fig. 13 demonstrates the process by which LLMs generate new examples given the ARC task and the corresponding ITP. LLMs generate multiple inputs that can form pairs with the output from one example of the task. This example, as mentioned earlier, is selected to serve as the basis for generation and is not included in ITP. If LLMs understood the specific rules corresponding to the ARC task given through ITP, the new example pairs generated by LLMs would be suitable as examples of the task.
@@
Before analyzing the experimental results, it was necessary to redefine the evaluation metric because the focus shifted from solving ARC tasks to generating valid examples. As previously explained, for a given example of a particular ARC task, we generated valid inputs that could be paired with the corresponding output. To successfully generate these inputs, the LLM must understand the specific rules of the given ARC task through its ITP and apply those understood rules to the output to create valid inputs. In this experiment, we evaluated whether all generated inputs were valid for each ARC task. This evaluation metric assesses both the LLM’s understanding of the correct rules and its ability to generate valid examples based on those rules. Consequently, this experiment systematically evaluates the LLMs’ capability to generate logical and valid demo pairs, enhancing our understanding of their ability to create new representations.
@@
Based on 160 ARC tasks classified by ConceptARC, we evaluated the validity of a total of 2,913 generated examples. The average valid generation ratio was approximately 17.1%, with the remaining examples deemed invalid. As previously mentioned, the validity of these results was determined by human judgment, assessing whether the generated tasks adhered to the given rules. The results in Table LABEL:tbl:productivity/generation_validity show that LLMs exhibit some capability in generating examples that align with the specified rules. However, due to weak criteria for determining validity, there is a limitation: even if an infinite number of results can be generated, they cannot be reliably used without post-processing the data.
3.3.3. Case Study: Invalid Production
An analysis of the generated inputs was conducted to investigate the reasons behind LLMs’ inability to produce valid inputs for ARC tasks. Two major limitations were observed when LLMs generated new ARC tasks. First, LLMs tended to simply copy inputs rather than infer meaningful rules from given example pairs. As shown in Fig. 15, this occurred repeatedly despite attempts to prevent it through prompts. Second, LLMs failed to properly consider the steps needed to generate inputs from given outputs. This frequently resulted in the creation of examples that could not be solved by the specific rules of the task. For instance, in cases where all vertices of a square were erased in the input, it became impossible to determine the color of the vertices, making it infeasible to infer the given output. These limitations suggest that LLMs lack an understanding of the semantics applicable in ARC tasks and the ability to compose these semantics according to constraints.
@@ 3.3.4. Conclusion
In Section 3.3, we conducted experiments to confirm the productivity of LLMs by assessing whether they can understand given ARC tasks in abstracted representations and generate valid new examples based on abstracted rules. Although it is known that LLMs have great strengths in creating creative works, our experimental results reveal that LLMs are weak in understanding rules and producing creations that adhere to those rules. Moreover, the observed limitations highlight a critical gap in LLMs’ ability to engage in higher-level reasoning and abstraction, which are essential for successfully solving ARC tasks that require an understanding of underlying principles rather than surface patterns. These results suggest that when LLMs generate outputs, they tend to mimic human-created results rather than truly understanding and applying rules. This makes it difficult for LLMs to reach the level of generation that humans can achieve.
Similarly, previous studies have shown similar results in measuring the productivity of AI models. Researchers tested how well pre-LLM models generalize to novel command combinations (Lake and Baroni, 2018; van der Velde et al., 2004). Their findings revealed strong performance on trained data but weaknesses in generating responses to unseen commands. Some researchers argued that LLMs struggle with generation under complex constraints and proposed improved models to address this issue (Koh et al., 2024; Liu et al., 2024). They propose novel frameworks to enhance LLMs for generating desired outputs when complex constraints are introduced, rather than relying solely on the base models. These researches share similarities with our study, which encountered difficulties in augmenting valid ARC tasks based on complex rules.
@@ 4.1.1. Abstract Knowledge
To solve ARC, the first challenge is the ability to extract the implicit information contained within ARC. Xu et al. argued that object-based representation is crucial for solving ARC and proposed ARGA (Xu et al., 2023a), which converts given example grids into graphs. Their follow-up study (Xu et al., 2023b) involved LLM solving ARC tasks using information obtained from ARGA and showed notable performance for object-based ARC tasks. However, these studies have a fundamental weakness in that they cannot be applied to ARC tasks without objects. Since only about 40% of ARC tasks contain object concepts (Xu et al., 2023a), this method cannot be applied to more than half of the tasks. Wang et al. on the other hand, improved the abstraction ability of LLMs to some extent with a graph-form dataset consisting of 221K textual descriptions, called AbsPyramid (Wang et al., 2023b), and also proposed a framework called AbsInstruct (Wang et al., 2024a) utilizing this dataset. Attempting to structure sentences can be an effective abstraction method for natural language, but its effectiveness cannot be seen in tasks that do not contain sentences.
@@ 4.1.2. Reasoning
Another challenge for LLMs in the context of ARC is the vast search space. One method gaining attention to address this is to enable LLMs to generate DSLs themselves. Rajani et al. introduced CAGE (Rajani et al., 2019), which prompts LLMs to generate explanations before generating answers. Subsequently, Wang et al. (Wang et al., 2024b) reported improved results by having LLMs generate DSLs based on hypotheses they set themselves. Additionally, active research is underway on prompting techniques applying algorithmic approaches. Zhou et al. (Zhou et al., 2022) demonstrated enhanced inference performance in LLMs by applying in-context learning. Follow-up research is actively being conducted following CoT and ToT. For example, CoT-SC (Wang et al., 2023c) is a study that selects results through voting from multiple instances of CoT, GoT (Besta et al., 2023) secures flexibility by enabling the generation of graph-like thought nodes, and XoT (Ding et al., 2023) uses the thought tree while Monte Carlo tree search and refines the tree with reinforcement learning. However, these attempts are closer to additional learning for LLMs, and more researches are needed to ascertain whether fundamental improvements in LLMs’ reasoning abilities are achievable.
@@ 4.3.3. Adding Evaluation Methods to Compare Task-Solving Processes with Human Approaches
4.4. Recent Research Trends on the Reasoning Abilities of LLMs
In this paper, we utilized the ARC to evaluate and enhance the reasoning capabilities of LLMs. ARC serves as a crucial benchmark for testing AI models’ ability to perform human-like reasoning, and it plays a central role in our research. Beyond ARC, various other datasets can be leveraged to strengthen LLMs’ reasoning abilities further. For instance, the DROP (Discrete Reasoning Over Paragraphs) dataset (Dua et al., 2019) evaluates a model’s capability to perform logical reasoning over complex text and synthesize information from multiple sources to generate accurate answers. Similarly, CommonsenseQA (Talmor et al., 2019) supports training models to perform commonsense reasoning, enabling them to make logical decisions in everyday scenarios. The BoolQ dataset (Clark et al., 2019) requires models to use logical reasoning to answer true/false questions, while the GSM8K dataset (Cobbe et al., 2021) focuses on enhancing step-by-step logical and mathematical reasoning skills through high-quality problem-solving tasks. Together, these datasets, alongside ARC, provide invaluable resources for systematically enhancing the diverse reasoning capabilities of LLMs.
However, recent studies indicate that despite their proficiency in language-based tasks, LLMs still exhibit significant limitations in their reasoning abilities. Carvalho et al. (de Carvalho et al., 2024) explored the reasoning capabilities of models like GPT-3.5 and GPT-4 in non-linguistic tasks requiring strategic thinking and spatial reasoning tasks. They found that while LLMs perform well on basic language tasks, they struggle with reasoning and decision-making in tasks beyond their training data, demonstrating limited cognitive flexibility and generalization. Similarly, Gendron et al. (Gendron et al., 2024) critically examined the abstract reasoning skills of LLMs, revealing poor performance on tasks requiring the identification and application of general patterns from only a small number of examples, specifically designed to test broad generalization abilities. These studies collectively highlight that current LLMs, though advanced in linguistic tasks, are still far from achieving robust reasoning abilities across diverse domains.
To build on this foundation, several advanced approaches have been developed to further enhance LLMs’ reasoning capabilities. Reinforcement learning with human feedback (Christiano et al., 2017) allows models to learn and refine their reasoning processes through feedback during training. The Chain of Thought (CoT) prompting technique (Wei et al., 2022) facilitates multi-step reasoning by breaking down complex tasks into smaller, manageable steps. Reasoning-centric fine-tuning (Lewkowycz et al., 2022) focuses on improving performance in specific reasoning tasks by leveraging specialized datasets. Additionally, incorporating knowledge graphs during pre-training (Liu et al., 2020) enhances logical inference by integrating structured world knowledge. Finally, explainable AI techniques, such as self-reflective learning, empower LLMs to identify and correct their reasoning errors, thereby contributing to greater transparency and trustworthiness in AI (Biran and Cotton, 2017). Collectively, these approaches play a crucial role in advancing LLMs’ reasoning capabilities across various domains.
Moreover, recent research has introduced innovative approaches to further augment the reasoning capabilities of LLMs. For example, multimodal learning techniques allow LLMs to simultaneously process text and image data, thereby enhancing their ability to tackle complex reasoning tasks (Singh et al., 2022). Adaptive learning strategies that incorporate human feedback have also been shown to improve the models’ adaptability across a wide range of tasks (Ouyang et al., 2022). Additionally, integrating programming languages with LLMs has been proposed as a means to enhance logical problem-solving abilities (Gao et al., 2023). These cutting-edge studies significantly contribute to systematically strengthening the multidimensional reasoning capabilities of LLMs.
@@ 5. Conclusions
In this study, we addressed the limitation of existing research, which predominantly analyzed LLM’s inference ability from a deterministic perspective, by introducing the LoTH perspective to ensure a fair evaluation of the process as well. By comparing and analyzing arguments across various domains of inference, we confirmed that logical coherence, compositionality, and productivity are quantifiable components to evaluate inference ability. Next, we proposed three experiments using the ARC to quantitatively evaluate these three components from the LoTH perspective. The results showed that although current LLMs exhibit outstanding performance, they lack logical coherence, compositionality, and productivity in their processes, suggesting that they are closer to probabilistic mimicry rather than possessing autonomous reasoning abilities. Finally, we explored meaningful research directions for LLM to acquire reasoning abilities from the LoTH perspective, as well as alternative approaches beyond ARC. This attempt to quantitatively evaluate the inference process from the LoTH perspective through experimental methods represents a differentiated approach not present in previous research. It contributes by providing a new perspective on how to handle reasoning abilities in the field of computer science, going beyond LLMs.
@@ Acknowledgements.
This work was supported by the IITP (RS-2023-00216011, RS-2024-00445087, No. 2019-0-01842) and the NRF (RS-2024-00451162) grants funded by the Ministry of Science and ICT, Korea. Experiments are supported by the Accelerate Foundation Models Research Initiative, Microsoft.References
@@
- Allen-Zhu and Li (2023) Zeyuan Allen-Zhu and Yuanzhi Li. 2023. Physics of Language Models: Part 3.2, Knowledge Manipulation. arXiv:2309:14402 (2023).
@@
- Biran and Cotton (2017) Or Biran and Courtenay Cotton. 2017. Explanation and Justification in Machine Learning: A Survey. In IJCAI Workshop.
@@
- Christiano et al. (2017) Paul F Christiano, Jan Leike, Tom Brown, Miljan Martic, Shane Legg, and Dario Amodei. 2017. Deep Reinforcement Learning from Human Preferences. In NeurIPS.
- Clark et al. (2019) Christopher Clark, Kenton Lee, Ming-Wei Chang, Tom Kwiatkowski, Michael Collins, and Kristina Toutanova. 2019. BoolQ: Exploring the Surprising Difficulty of Natural Yes/No Questions. In NAACL.
- Cobbe et al. (2021) Karl Cobbe, Vineet Kosaraju, Mohammad Bavarian, Mark Chen, Heewoo Jun, Lukasz Kaiser, Matthias Plappert, Jerry Tworek, Jacob Hilton, Reiichiro Nakano, Christopher Hesse, and John Schulman. 2021. Training Verifiers to Solve Math Word Problems. arXiv:2110.14168 (2021).
@@
- de Carvalho et al. (2024) Gonçalo Hora de Carvalho, Robert Pollice, and Oscar Knap. 2024. Show, Don’t Tell: Evaluating Large Language Models Beyond Textual Understanding with ChildPlay. arXiv:2407.11068 (2024).
@@
- Dua et al. (2019) Dheeru Dua, Yizhong Wang, Pradeep Dasigi, Gabriel Stanovsky, Sameer Singh, and Matt Gardner. 2019. DROP: A Reading Comprehension Benchmark Requiring Discrete Reasoning Over Paragraphs. In NAACL.
@@
- Fodor and Pylyshyn (1988) Jerry A Fodor and Zenon W Pylyshyn. 1988. Connectionism and Cognitive Architecture: A Critical Analysis. Cognition 28, 1-2 (1988), 3–71.
- Gao et al. (2023) Luyu Gao, Aman Madaan, Shuyan Zhou, Uri Alon, Pengfei Liu, Yiming Yang, Jamie Callan, and Graham Neubig. 2023. PAL: Program-Aided Language Models. In ICML. 10764–10799.
@@
- Gendron et al. (2024) Gaël Gendron, Qiming Bao, Michael Witbrock, and Gillian Dobbie. 2024. Large Language Models are Not Strong Abstract Reasoners. IJCAI (2024).
@@
- Hodel (2024) Michael Hodel. 2024. Addressing the Abstraction and Reasoning Corpus via Procedural Example Generation. arXiv:2404.07353 (2024). https://github.com/michaelhodel/re-arc?tab=readme-ov-file
- Huang and Chang (2023) Jie Huang and Kevin Chen-Chuan Chang. 2023. Towards Reasoning in Large Language Models: A Survey. In ACL Findings.
- Hupkes et al. (2020) Dieuwke Hupkes, Verna Dankers, Mathijs Mul, and Elia Bruni. 2020. Compositionality Decomposed: How Do Neural Networks Generalise? Journal of Artificial Intelligence Research 67 (2020), 757–795.
@@
- Koh et al. (2024) Jing Yu Koh, Daniel Fried, and Russ R Salakhutdinov. 2024. Generating Images with Multimodal Language Models. In NeurIPS, Vol. 36.
- Lab42 (2024) Lab42. 2024. ARC-PRIZE competition. https://arcprize.org/
- Lake and Baroni (2018) Brenden Lake and Marco Baroni. 2018. Generalization without Systematicity: On the Compositional Skills of Sequence-to-Sequence Recurrent Networks. In ICML. PMLR, 2873–2882.
@@
- Lewkowycz et al. (2022) Aitor Lewkowycz, Anders Andreassen, David Dohan, Ethan Dyer, Henryk Michalewski, Vinay Ramasesh, Ambrose Slone, Cem Anil, Imanol Schlag, Theo Gutman-Solo, Yuhuai Wu, Behnam Neyshabur, Guy Gur-Ari, and Vedant Misra. 2022. Solving Quantitative Reasoning Problems with Language Models. In NeurIPS.
@@
- Liu et al. (2024) Mushui Liu, Yuhang Ma, Xinfeng Zhang, Yang Zhen, Zeng Zhao, Zhipeng Hu, Bai Liu, and Changjie Fan. 2024. LLM4GEN: Leveraging Semantic Representation of LLMs for Text-to-Image Generation. arXiv:2407.00737 (2024).
- Liu et al. (2020) Weijie Liu, Peng Zhou, Zhe Zhao, Zhiruo Wang, Qi Ju, Haotang Deng, and Ping Wang. 2020. K-BERT: Enabling Language Representation with Knowledge Graphs. In AAAI.
@@
- Ma et al. (2024) Teli Ma, Rong Li, and Junwei Liang. 2024. An Examination of the Compositionality of Large Generative Vision-Language Models. In NAACL.
@@
- Ouyang et al. (2022) Long Ouyang, Jeffrey Wu, Xu Jiang, Diogo Almeida, Carroll Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, John Schulman, Jacob Hilton, Fraser Kelton, Luke Miller, Maddie Simense, Amanda Askell, Peter Welinder, Paul Christiano, Jan Leike, and Ryan Lowe. 2022. Training Language Models to Follow Instructions with Human Feedback. In NeurIPS.
@@
- Singh et al. (2022) Amanpreet Singh, Ronghang Hu, Vedanuj Goswami, Guillaume Couairon, Wojciech Galuba, Marcus Rohrbach, and Douwe Kiela. 2022. FLAVA: A Foundational Language and Vision Alignment Model. In CVPR.
- Sinha et al. (2024) Sania Sinha, Tanawan Premsri, and Parisa Kordjamshidi. 2024. A Survey on Compositional Learning of AI Models: Theoretical and Experimental Practices. arXiv:2406.08787 (2024).
- Talmor et al. (2019) Alon Talmor, Jonathan Herzig, Nicholas Lourie, and Jonathan Berant. 2019. CommonsenseQA: A Question Answering Challenge Targeting Commonsense Knowledge. In NAACL.
@@
- Tyen et al. (2024) Gladys Tyen, Hassan Mansoor, Victor Cărbune, Yuanzhu Peter Chen, and Tony Mak. 2024. LLMs cannot find reasoning errors, but can correct them given the error location. In ACL Findings.
@@
- van der Velde et al. (2004) Frank van der Velde, Gwendid T van der Voort van der Kleij, and Marc de Kamps. 2004. Lack of Combinatorial Productivity in Language Processing with Simple Recurrent Networks. Connection Science 16, 1 (2004), 21–46.
- Wang et al. (2023a) Boshi Wang, Sewon Min, Xiang Deng, Jiaming Shen, You Wu, Luke Zettlemoyer, and Huan Sun. 2023a. Towards Understanding Chain-of-Thought Prompting: An Empirical Study of What Matters. In ACL.
@@
- Wang et al. (2023c) Xuezhi Wang, Jason Wei, Dale Schuurmans, Quoc Le, Ed Chi, Sharan Narang, Aakanksha Chowdhery, and Denny Zhou. 2023c. Self-Consistency Improves Chain of Thought Reasoning in Language Models. In ICLR.
@@
- Wang et al. (2023b) Zhaowei Wang, Haochen Shi, Weiqi Wang, Tianqing Fang, Hongming Zhang, Sehyun Choi, Xin Liu, and Yangqiu Song. 2023b. AbsPyramid: Benchmarking the Abstraction Ability of Language Models with a Unified Entailment Graph. arXiv:2311.09174 (2023).
@@
- Yang et al. (2024) Haoran Yang, Hongyuan Lu, Wai Lam, and Deng Cai. 2024. Exploring Compositional Generalization of Large Language Models. In NAACL Workshop.
@@
- Zhang et al. (2024b) Jinghan Zhang, Xiting Wang, Weijieying Ren, Lu Jiang, Dongjie Wang, and Kunpeng Liu. 2024b. RATT: A Thought Structure for Coherent and Correct LLM Reasoning. arXiv:2406.02746 (2024).
- Zhang et al. (2024a) Wenqi Zhang, Yongliang Shen, Linjuan Wu, Qiuying Peng, Jun Wang, Yueting Zhuang, and Weiming Lu. 2024a. Self-Contrast: Better Reflection Through Inconsistent Solving Perspectives. In ACL.